2. Given the function g(t) = 2t^2– t – 1, which of the following equals g(2 + 3 )?
Answer by stanbon(75887)

Given the function g(t) = 2t^2– t – 1, which of the following equals g(2 + 3 )?
------
g(2+3) = g(5) = 2*5^2-5-1 = 25-6 = 19
-------------
Cheers,
Stan H.
-------------